OpenStudy (anonymous):

10 + h/3 = 1 ? Solve for h and show steps please

OpenStudy (anonymous):

subtract 10 on both sides, multiply 3 on both sides

OpenStudy (anonymous):

so, 1-10=9 9*3=27

OpenStudy (anonymous):

27 is ur answer

sammixboo (sammixboo):

First you have to subtract 10 from both sides \(\sf \large \color{magenta}{10~-~10~+~h/3~=~1~-~10}\)

sammixboo (sammixboo):

So what is \(\sf \large \color{magenta}{10~-~10}\) and \(\sf \large \color{magenta}{1~-~10}\)?

sammixboo (sammixboo):

\(\sf \large \color{magenta}{10~-~10~=~0}\) and \(\sf \large \color{magenta}{1~-~10~=~-9}\) So now our problem looks like this \(\sf \large \color{magenta}{h/3~=~-9}\)

sammixboo (sammixboo):

So now we take 3 and mltiply it on both sides \(\sf \large \color{magenta}{h/3~x~3~=~-9~x~3}\) Since \(\sf \large \color{magenta}{h/3~x~3}\) gets canceled out it is just h \(\sf \large \color{magenta}{h~=~-9~x~3}\) \(\sf \large \color{magenta}{h~=~-27}\)